OpenStudy (anonymous):

trains A and B are traveling in the same direction on a parallel track. train A is traveling at 40 miles per hours and train B is traveling at 50 miles per hour. train a passes a staion at 7:25 pm. if train B passes the same station at 7:37 pm at what time will train B catch up to train A?

OpenStudy (anonymous):

i came up with diff answers so im not for sure i got 8:25

OpenStudy (anonymous):

725-737= is 12 min difference in the two

OpenStudy (anonymous):

thats 1/5 of an hour

OpenStudy (a_clan):

let t1hrs be the time for train A (speed 40 mph) to reach 'catching' point and t2 hrs for train B (speed 50 mph). Since A started 12 minute earlier t1=t2+12 now since both travelled same distance in reaching the catching point t1 * s1 = t2 * s2 (t2+12)*40 = t2*50 t2= 48 hrs Ans: 48 hrs after 7:37 pm i.e. 2 days after, at the same time.

OpenStudy (anonymous):

mult. 1/5*40=8 miles 40-8=32 miles difference

OpenStudy (anonymous):

so exactly what is the ans

OpenStudy (anonymous):

theres 10 mph difference in the two

OpenStudy (anonymous):

so tena32 at what time with train b catch up with train a

OpenStudy (anonymous):

im getting 8:25

OpenStudy (a_clan):

what approach did you follow?

OpenStudy (anonymous):

let me type it

OpenStudy (a_clan):

ok

OpenStudy (anonymous):

2:37-2:25= 12 min diff in the two trains which is 1/5 hour the difference in speed is 10 (50-40) 1/5*40=8

OpenStudy (anonymous):

tena32 is it 8:25 am

OpenStudy (a_clan):

it is not that simple

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

it is 8.25pm

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

no this problem is horrible ive had my niece work it and she is in honors in college and she had troublr

OpenStudy (anonymous):

i still have like 3 more word problems to go lol

OpenStudy (a_clan):

@tan1992: what is the solution?

OpenStudy (anonymous):

ok i try my best to explain

OpenStudy (anonymous):

yall gonna love me by the time im done lol

OpenStudy (a_clan):

@tena32: It is not that horrible as well. Just that I have been out of touch

myininaya (myininaya):

heu o got 8:25 too

myininaya (myininaya):

i will write it out nicely and post it k?

OpenStudy (anonymous):

train b is 10 mile faster than train a per hour train a reach station at 7.25 while train b reach 7.37 which mean when train reach the station train a is moving at 40 mile per hour for 12 min= 8 MILE since b can catch up a 10 mph ,then it take 48 MIN to catch up 8 mile 7.37pm+48min=8.25

OpenStudy (anonymous):

thats it tan. i was on the right track no pun intended

OpenStudy (a_clan):

since b can catch up a 10 mph ,then it take 48 MIN to catch up 8 mile @tan1992: Please Explain?

OpenStudy (anonymous):

mean train b faster than train a 10 mile per hour we use 8 mile dive by 10mph we will get 0.8 which mean train b need 0.8 hour to catch up train a

OpenStudy (anonymous):

*devide

myininaya (myininaya):

i got excited so i hope you can read it

myininaya (myininaya):

i made equation for each train and i found when train B will catch up to train A by setting these equations equal

OpenStudy (a_clan):

@myininaya: which co-ordinate system are you using? position - time ?

myininaya (myininaya):

rectangular

myininaya (myininaya):

cartesian

myininaya (myininaya):

time is x axis position is y right!

myininaya (myininaya):

does it make sense? do i need to explain anything?

OpenStudy (a_clan):

@Myininnaya: how did you arrive at the equations of line

OpenStudy (a_clan):

@tan1992: I still am not sure if we could use the difference in speed of both the trains i.e. 10 mph to find anything in this problem unless we are using relative motion.

myininaya (myininaya):

train A time=t t0=0 t1=0+40 t2=40+40 ... tx=40x train B time=t t0=0 t1=0+50 t2=50+50 .... tx=50x but they want to know from the when they hit that station so train A has position y=40x+445/60 445/60 is 7:25 in minutes train B has position y=50x+457/60

myininaya (myininaya):

since 457/60 minutes=7:37

myininaya (myininaya):

50x+457/60=40x+445/60 gives x=101/12=8 25/60 =8:25

myininaya (myininaya):

tx means at time x

OpenStudy (a_clan):

@myininaya:: y denotes a position. right? So for A, y= 40x + ---- (you used 445/60 min which is a time quantity) can you add to different units to obtain y( a position)?

myininaya (myininaya):

oops by the way those lines above were meant to be y=40x-445/60 and y=50x-457/60 sorry for type-0 aclan give me a second it might help if i explain both parts together

OpenStudy (a_clan):

@myininaya:: And does the slope of position-time graph give speed? As I can recall, it was distance - time graph.

myininaya (myininaya):

my explanation is sort of bad above but... okay we aren't looking at the station right now okay? train A at time 0 hour, the position is 0 at time 1 hour, the poisiton is 40 at time 2 hour, the position is 40+40=40(2) at time 3 hour, the position is 40+40+40=40(3) ... at time x hour, the position is 40x now lets keep the station in mind now so if we draw the position 0 is at hour 7 25/60 then we are on the x-axis so we have a line y=mx+b we know a point on the line (7 25/60,0) and m=40 0=40*445/60+b so the y intercept is -40*445/60 so the position of train A is y=4x-40*445/60 the position of the train is only changing due to speed of the train and they give us this in the problem the train is traveling 40 miles per hour

myininaya (myininaya):

similiary we can do the same process for train B

myininaya (myininaya):

the position is y=speed*time-blah

myininaya (myininaya):

a clan do you understand the train is moving due to the speed right?

myininaya (myininaya):

at time 0, we are at position 0 at time 1, we are at poistion 0+40 ... time x, we are at position 40x do you not undestand this part?

myininaya (myininaya):

x is time

OpenStudy (a_clan):

I understand

myininaya (myininaya):

the train is moving 40 miles every hour

OpenStudy (a_clan):

so final answer would be----?

myininaya (myininaya):

lol 8:25 same as what i got along itme ago